# Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study

### v2

@@ Abstract
Effective toxic content detection relies heavily on high-quality and diverse data, which serves as the foundation for robust content moderation models. This study explores the potential of open-source LLMs for harmful data synthesis, utilizing prompt engineering and fine-tuning techniques to enhance data quality and diversity. In a two-stage evaluation, we first examine the capabilities of six open-source LLMs in generating harmful data across multiple datasets using prompt engineering. In the second stage, we fine-tune these models to improve data generation while addressing challenges such as hallucination, data duplication, and overfitting. Our findings reveal that Mistral excels in generating high-quality and diverse harmful data with minimal hallucination. Furthermore, fine-tuning enhances data quality, offering scalable and cost-effective solutions for augmenting datasets for specific toxic content detection tasks. These results emphasize the significance of data synthesis in building robust, standalone detection models and highlight the potential of open-source LLMs to advance smaller downstream content moderation systems. We implemented this approach in real-world industrial settings, demonstrating the feasibility and efficiency of fine-tuned open-source LLMs for harmful data synthesis.
@@ 1 Introduction
Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
In toxic content detection, (1) labeling real-world datasets is both expensive and labor-intensive, with the substantial volume of data required further compounding the challenge—our costs have reached millions of dollars. (2) Ensuring dataset diversity and quality remains a persistent issue, often hindering the development of robust detection models. (3) The inherent ambiguity and subjectivity of the term "harmful"—which varies across topics and annotator perspectives—adds another layer of complexity. Data synthesis presents a promising approach to mitigating these challenges. Moreover, many product applications demand real-time toxic content detection with stringent latency requirements (e.g., latency $\leq$ 0.01 seconds). Directly using SOTA Podolak et al. (2024); Fillies and Paschke (2024) LLM-based approaches might not reliably meet these constraints, making LLM-generated data synthesis for training smaller, more robust classification models a practical solution.
This study explores the use of prompt engineering and fine-tuning techniques with open-source LLMs to improve harmful data synthesis. Existing methods, such as ToxiGen Hartvigsen et al. (2022) and Toxicraft Hui et al. (2024) provide useful frameworks but rely on GPT-based frameworks for data generation, which face issues of cost and scalability. For example, ToxiGen employs GPT3 within the ALICE framework, producing high-quality datasets but at a high computational expense. These limitations highlight the need for alternative solutions. Our study addresses these challenges by investigating how open-source LLMs can facilitate harmful data synthesis. Open-source models offer comparable effectiveness to proprietary systems while reducing computational costs and enhancing explainability Ding et al. (2024). Many products can benefit from standalone data synthesis modules that generate training data for domain-specific and time-sensitive content safety applications, making open-source LLMs a scalable and cost-effective choice.
Our experimental approach includes a comprehensive evaluation of six open-source LLMs: Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024), explored across diverse prompt engineering strategies Schulhoff et al. (2024); Li (2023) and fine-tuning configurations. The research is divided into two key stages: In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al. (2024). However, we observed that relying solely on prompt engineering often results in suboptimal outputs due to the model’s inherent safety alignments, which limit harmful data generation. Consequently, we conducted a second stage of experiments involving fine-tuning. In the second stage, we fine-tuned the LLMs using proprietary datasets and explored various configurations, such as epoch settings and data mixing strategies. Our results demonstrate that fine-tuning effectively mitigates issues like hallucination in LLMs and enhances the quality of harmful data generation. However, challenges such as data duplication and overfitting were identified, emphasizing the need for a careful balance between training epochs and dataset diversity.
In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas tailored for industrial applications of toxic content moderation systems. These efforts provide key insights into optimizing open-source LLMs for harmful data synthesis, enhancing both the robustness and accuracy of toxic content detection systems. Additionally, this study validates the practicality and efficiency of fine-tuned open-source LLMs in real-world industrial applications. Our main contributions are summarized as follows:
@@
  • Our study is one of the first (if not the first) to apply prompt engineering and fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
@@ 2 Methodology
Our approach utilizes prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data synthesis, thereby improving toxic content detection systems. The methodology comprises two key stages: prompt engineering and fine-tuning, as illustrated in Figure 1.
@@ 2.1 Prompt Engineering
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s inherent safety alignments, which are explicitly designed to prevent harmful content generation. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked sufficient quality and diversity, hindering their utility in effective toxic content detection. To overcome these challenges, we experimented with various configurations, such as altering few-shot examples and adjusting requirement specificity. However, these adjustments yielded limited success in bypassing the internal safety mechanisms. After testing, the final prompt template are showed in Appendix A. Our findings showed that relying solely on prompt engineering was inadequate for our objectives. This realization prompted the integration of a second stage—fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
2.2 Fine-Tuning
Based on the results from prompt engineering, we hypothesized that fine-tuning would yield a more effective solution. To achieve this, we employed LoRA methods Hu et al. (2022) for fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
@@ 2.2 Fine-Tuning
After fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 4.
@@ 2.3 Algorithm
Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training. Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
@@ 3.1 Datasets
Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicate in bold.
@@ 3.2 Models
Stage 1: Prompt Engineering
n the first stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al. (2023), LLaMA2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024). Detailed results and analysis for Stage 1 are provided in Section 4.1. Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the fine-tuning stage.
Stage 2: Fine-Tuning
In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al. (2022). After comparing latency and cost we decided to use a smaller model Mixtral for further exploration. Detailed results and analysis for Stage 2 are provided in Section 4.2.
@@ 3.3 Implementation
The fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al. (2024). We employed a Supervised Fine-Tuning (SFT) methodology with a Low-Rank Adaptation (LoRA) setup Hu et al. (2022). Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B. The results of this fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 4.
@@
In our production environment, meeting strict inference time requirements is critical. Among the evaluated models, the Multilayer Perceptron (MLP) Gardner and Dorling (1998) stood out for its ability to fulfill these demands. For downstream model evaluation, in this paper we utilized a MLP classifier with two hidden layers, each comprising 600 neurons.
4 Result
This section presents the results of our study, showcasing the effectiveness of both prompt engineering and fine-tuning strategies in optimizing open-source LLMs for harmful data synthesis. Our evaluation employed a two-stage approach: Stage 1 focused on prompt engineering to assess the initial performance of various models, while Stage 2 involved fine-tuning the selected models to further enhance data quality and diversity.
@@ 4.1 Stage 1: Prompt Engineering
The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data synthesis using prompt engineering. The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content. Table 2 presents the performance comparison of the six different models evaluated in Stage 1. The success rates for political and hate content are reported, along with the results of human evaluation.
The success rate columns in Table 2 represent the percentage of valid positive samples generated by each model for political and hate content. The values indicate the proportion of samples meeting the minimum requirements (60% of the total samples generated). The human evaluation columns show whether the generated data was subject to human evaluation. A green checkmark () signifies that human annotators evaluated the quality of the generated data and found a high Fleiss’ kappa score, indicating strong inter-rater agreement. This also means that the generated samples met the minimum quality criteria. Based on the evaluation, Mistral and Vicuna demonstrated the best performance with prompt engineering. These two models were selected for the fine-tuning stage due to their superior results in generating high-quality and diverse data entries.
@@
Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best result are indicate in bold.
4.2 Stage 2: Fine-tuning
In the second stage, we began by fine-tuning the models exclusively on individual datasets and evaluating their performance on the same datasets. The results of this approach are detailed in Table 4. Inspired by Hui et al. (2024), Yu et al. (2024), and Qi et al. (2023), we subsequently adopted a data mixing strategy that combined multiple datasets for fine-tuning. The results of this multi-dataset fine-tuning are presented in Table 3. Additionally, we conducted an ablation study to examine the impact of different design choices, detailed in Appendix D. The "Mixture" versions represent fine-tuning on a mix of hate, violence, and sex datasets, while the "Mix_GPT" versions incorporate GPT-generated positive samples into the mixed data. For all fine-tuning experiments, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template. Our findings reveal that fine-tuning significantly enhances the performance of LLMs in generating harmful data. The data mixing strategy helps in mitigating issues like data duplication and overfitting. The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications. One surprising observation was the notable improvement in model performance when using mixed data, suggesting that diverse training data enables better generalization and more reliable outputs. The Mixture models trained for Epoch 3 and Epoch 5 showed particularly strong results, emphasizing the importance of sufficient training epochs for optimal performance.

### v3

# ToxiLab: How Well Do Open-Source LLMs Generate Synthetic Toxicity Data?

@@ Abstract
Effective toxic content detection relies heavily on high-quality and diverse data, which serve as the foundation for robust content moderation models. Synthetic data has become a common approach for training models across various NLP tasks. However, its effectiveness remains uncertain for highly subjective tasks like hate speech detection, with previous research yielding mixed results. This study explores the potential of open-source LLMs for harmful data synthesis, utilizing controlled prompting and supervised fine-tuning techniques to enhance data quality and diversity. We systematically evaluated six open source LLMs on five datasets, assessing their ability to generate diverse, high-quality harmful data while minimizing hallucination and duplication. Our results show that Mistral consistently outperforms other open models, and supervised fine-tuning significantly enhances data reliability and diversity. We further analyze the trade-offs between prompt-based vs. fine-tuned toxic data synthesis, discuss real-world deployment challenges, and highlight ethical considerations. Our findings demonstrate that fine-tuned open source LLMs provide scalable and cost-effective solutions to augment toxic content detection datasets, paving the way for more accessible and transparent content moderation tools.
@@ 1 Introduction
Toxic content detection requires large-scale labeled datasets, but manual annotation is costly and labor-intensive. The sheer volume of required data has driven costs into the millions. Maintaining dataset diversity and quality remains a persistent challenge, often hindering the development of robust detection models. Additionally, the definition of "harmful" content is inherently subjective, varying across topics and annotator perspectives, adding complexity to dataset curation Casula and Tonelli (2023). Synthetic data generation Zhang et al. (2022); Ye et al. (2022b) presents a scalable, cost-effective solution to these challenges. Training models for various NLP tasks using synthetic data has become increasingly common. However, previous research has reported mixed results regarding its effectiveness in highly subjective tasks such as hate speech detection Casula et al. (2024). Many existing synthetic data methods for hate speech detection rely on simple prompt-based rewrites or sentence reordering, often failing to capture nuanced toxicity patterns. Furthermore, most of the best-performing methods have utilized proprietary GPT-series models, limiting scalability and accessibility.Existing methods such as ToxiGen Hartvigsen et al. (2022) and Toxicraft Hui et al. (2024) rely on proprietary GPT-based models and prompt engineering, incurring high computational costs and scalability limitations. These constraints underscore the need for open-source alternatives, motivating this study to investigate three key research questions:
  • How effectively can open-source LLMs generate synthetic toxic data compared to proprietary models?
  • What limitations arise from using prompt-based synthetic data generation for hate speech detection, and how can supervised fine-tuning improve performance?
  • How do different synthetic data generation strategies impact the robustness of toxic content detection models?
To address these questions, we systematically evaluate 6 open-source LLMs on 5 datasets. Our evaluation approach consists of two key stages:
  1. 1.
    Prompt Engineering – We compare LLMs, design structured prompts, and experiment with few-shot techniques to balance data quality and diversity Ding et al. (2024). However, we observe that prompt engineering alone is insufficient, as inherent safety alignments in LLMs restrict harmful content generation
  2. 2.
    supervised fine-tuning – We fine-tune LLMs using proprietary datasets, exploring configurations such as epoch settings and data mixing strategies. supervised fine-tuning improves data reliability, mitigates hallucination, and enhances dataset diversity, though challenges such as data duplication and overfitting persist.
In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas. Given that generated data is increasingly used even in sensitive applications Ghanadian et al. (2024), it becomes crucial for the NLP community to critically examine the impact of synthetic data—including its ethical risks—in a manner similar to discussions in other research communities Liu et al. (2024); Ai et al. (2024). Our work serves as an initial contribution in this direction, providing insights into the practical challenges of synthetic toxic data generation and its implications for real-world content moderation systems. By exploring the effectiveness of prompt engineering and supervised fine-tuning in open-source LLMs, we offer guidance for the responsible deployment of synthetic data in NLP applications. In summary, our main contributions are as follows:
@@
  • Our study is one of the first (if not the first) to apply prompt engineering and supervised fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
@@
Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for supervised fine-tuning (stage 2).
2 Related Work
2.1 Hate Speech Detection
Early studies established classifiers to detect harmful information using neural network models Zhang et al. (2018) or word embedding methods Kshirsagar et al. (2018). In recent years, models based on the Transformer architecture have demonstrated remarkable capabilities, prompting researchers to explore further. Rajput et al. (2021) conducted research on the ETHOS hate speech detection dataset, comparing classifiers’ performance in hate speech detection by replacing or integrating word embeddings (fastText, GloVe, or FT + GV) with BERT embeddings. Aluru et al. (2020) contrasted simple models (such as LASER embeddings with logistic regression) and BERT models in scenarios with scarce and abundant linguistic resources. Lin et al. (2024) generated explanations through multimodal debates between LLMs, enhancing the transparency and explainability of harmful meme detection. da Silva Oliveira et al. (2024) validated the effectiveness of ChatGPT in identifying harmful Spanish-language speech.
2.2 Data Synthesis
Traditional data synthesis methods have employed various techniques, ranging from synonym replacement to token-level manipulations, as exemplified by Wei and Zou (2019). While these methods provide some level of augmentation, they often lack contextual richness and diversity. The introduction of translation-based approaches (Fadaee et al., 2017) and masked language modeling (Kumar et al., 2020) improved semantic consistency, yet they still struggle to generate high-quality synthetic data necessary for complex tasks such as harmful content detection. Recent advancements in zero-shot data generation, such as ZEROGEN (Ye et al., 2022c), SuperGen (Meng et al., 2022), and PROGEN (Ye et al., 2022a), have explored methods for dataset synthesis without requiring extensive labeled examples. However, these frameworks often suffer from issues related to low information density and redundancy, limiting their effectiveness in generating diverse and contextually relevant samples.
@@ 3 Methodology
Our methodology systematically evaluates open-source LLMs for harmful data synthesis through a two-stage approach: prompt engineering and supervised fine-tuning., as illustrated in Figure 1.
@@ 3.1 Prompt Engineering
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s inherent safety alignments, which are explicitly designed to prevent harmful content generation. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked sufficient quality and diversity, hindering their utility in effective toxic content detection. To overcome these challenges, we experimented with various configurations, such as altering few-shot examples and adjusting requirement specificity. However, these adjustments yielded limited success in bypassing the internal safety mechanisms. After testing, the final prompt template are showed in Appendix A. Our findings showed that relying solely on prompt engineering was inadequate for our objectives. This realization prompted the integration of a second stage—supervised fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
3.2 supervised fine-tuning
Based on the results from prompt engineering, we hypothesized that supervised fine-tuning would yield a more effective solution. To achieve this, we employed LoRA methods Hu et al. (2022) for supervised fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The supervised fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
@@
After supervised fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 5.
@@ 3.3 Algorithm
Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for supervised fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training. Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
@@ 4.1 Datasets
Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicated in bold.
@@ 4.2 Models
Prompt Engineering
In this stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al. (2023), LLaMA2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024) and our baseline: GPT4 model Brown et al. (2020). Detailed results and analysis for Stage 1 are provided in Section 5.1. Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the supervised fine-tuning stage.
Supervised Fine-tuning
In supervised fine-tuning, we fine-tuned the Mistral Jiang et al. (2023) and Vicuna Chiang et al. (2023) models using the LoRA method Hu et al. (2022). After comparing latency and cost, we experimented and reported the better performance model Mixtral for further exploration. Detailed results and analysis for Stage 2 are provided in Section 5.2.
@@ 4.3 Implementation
The supervised fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al. (2024). We employed a supervised fine-tuning (SFT) methodology with a Low-Rank Adaptation (LoRA) setup Hu et al. (2022). Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B. The results of this supervised fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 5.
@@
We utilized a Multilayer Perceptron (MLP) Gardner and Dorling (1998) for downstream model evaluation. The MLP classifier consisted of two hidden layers, each comprising 600 neurons, with ReLU activation functions and a softmax output layer for classification. The model was trained using the Adam optimizer. We conducted multiple training runs with different random seeds to ensure stability and robustness in evaluation. Given its simple structure and strong interpretability, MLP provides a clear and reliable baseline for assessing the impact of synthetic data on model performance across various datasets and experimental conditions.
5 Results and Analysis
We present the results of our study, structured around the three research questions. Our two-stage evaluation—prompt engineering and supervised fine-tuning—allowed us to assess the effectiveness of open-source LLMs in synthetic toxic data generation, examine the limitations of prompt-based methods, and analyze the impact of different generation strategies on model robustness.
@@ 5.1 Stage 1: Prompt Engineering
The first stage of our evaluation assessed how well open-source LLMs could generate synthetic data using prompt engineering. The primary objective was to evaluate their ability to generate high-quality augmented data, particularly for political and hate speech content. Table 2 presents a performance comparison of six open-source models, including their success rates for political and hate content as well as human evaluation results. The success rate columns in Table 2 indicate the percentage of valid positive samples generated by each model, where a minimum threshold of 60% of the total samples was required. The human evaluation columns confirm whether the generated data underwent human annotation. A green checkmark () signifies that annotators validated the data quality, achieving a high Fleiss’ kappa score, which indicates strong inter-rater agreement.
5.2 Stage 2: Supervised Fine-tuning
In the second stage, models were supervised fine-tuned on individual datasets and evaluated accordingly, with results detailed in Table 4. Inspired by Hui et al. (2024), Yu et al. (2024), and Qi et al. (2023), we then explored data mixing, incorporating multiple datasets for training. Table 3 presents results from this multi-dataset fine-tuning strategy. Further analysis, including an ablation study on different design choices, is provided in Figure  2 and Appendix D. The "Mixture" versions in Table 3 represent models trained on a blend of hate, violence, and sex-related datasets. The "Mix_GPT" versions extend this approach by integrating GPT-generated positive samples. Across all fine-tuning experiments, we maintained a balanced dataset of 3000 positive and 3000 negative samples, formatted using the Alpaca instruction template.
@@ 5.2 Stage 2: Supervised Fine-tuning
Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best results are indicated in bold..
Figure 2: Ablation Study on Data mixing
5.3 RQ1: How effectively can open-source LLMs generate synthetic toxic data compared to proprietary models?
We compared six open-source LLMs against a GPT-4 baseline using prompt-based generation. Mistral and Vicuna consistently achieved the highest success rates and human evaluation scores among open models (Table 2). However, even these models struggled to match GPT-4 in data diversity and nuance, often producing repetitive or formulaic outputs. Fine-tuning significantly reduced this gap. A fine-tuned Mistral model demonstrated improved F1 scores and accuracy across multiple datasets (Table 4). Although GPT-4 remained the strongest performer, fine-tuned Mistral approached its performance, demonstrating that open-source LLMs, when fine-tuned, can serve as viable alternatives for synthetic data generation.
5.4 RQ2: What limitations arise from using prompt-based synthetic data generation for hate speech detection, and how can fine-tuning improve performance?
Prompt-based generation exhibited several critical limitations. Safety alignments in models such as Gemma and Bloom led to frequent refusals to generate toxic content, reducing dataset completeness. Additionally, generated samples often lacked diversity, limiting their utility for training robust detection models. Another issue was over-simplicity, as models struggled to produce nuanced toxicity, instead generating overly blunt or obvious harmful statements. Fine-tuning alleviated these challenges. Fine-tuned models demonstrated improved data diversity, generating more varied and contextually rich toxic content (Table 3). They also exhibited higher data reliability, reducing hallucination rates and enhancing dataset quality. Furthermore, fine-tuning enabled models to generate more subtle and realistic toxic content, making them more applicable to real-world hate speech detection. However, fine-tuning also introduced challenges such as data duplication and overfitting, particularly in later training epochs (Table 4).
5.5 RQ3: How do different synthetic data generation strategies impact the robustness of toxic content detection models?
We analyzed how different synthetic data generation strategies influenced classification performance. Fine-tuned models consistently outperformed those relying solely on prompt-generated data, achieving higher F1 scores and accuracy across multiple datasets (Table 4). Models trained on blended datasets containing multiple categories of toxic content (e.g., hate speech, violence, and sexual content) demonstrated stronger generalization compared to models trained on a single category. However, extended fine-tuning introduced overfitting risks. Performance improvements plateaued beyond epoch 3, and some models experienced performance degradation at epoch 5 (Table 4), likely due to memorization effects. Ablation studies (Appendix D) confirmed that strategic fine-tuning with mixed data provides the best balance between data quality and generalization, underscoring the importance of careful dataset curation and tuning strategies.